# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 7.4. Proof of Part (a)
For $i<q^{n-1}$, write $i=\ell q^{n-2}+r$ with $\ell<q$ and $r<q^{n-2}$. We will prove by induction on $i$ that $u_{m}^{\ell+1}\mathcal{P}^{i}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$. For $i=0$, we have $\ell=0$ and $u_{m}\mathcal{P}^{0}(d_{j,m})=u_{m}d_{j,m}\in R_{n-1}$ with $\xi_{n-1}$-degree $1$. For $0<i<q^{m-1}$, we have $u_{m}\mathcal{P}^{i}(d_{j,m})=\mathcal{P}^{i}(u_{m}d_{j,m})\in R_{n-2}$, giving $\mathcal{P}^{i}(d_{j,m})\in R_{n-2}$. For $i\geq q^{m-1}$, using the Cartan identity gives
@@
Since $u_{m}d_{j,m}\in R_{n-1}$ with $\xi_{n-1}$-degree $1$, we see that $\mathcal{P}^{i}(u_{m}d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$. For $k>0$, write $i-k=\ell^{\prime}q^{n-2}+r^{\prime}$ with $\ell^{\prime}\leq\ell$ and $r^{\prime}<q^{n-2}$. Since $i-k<i$, the induction hypothesis gives $u_{m}^{\ell^{\prime}+1}\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell^{\prime}+1$. If $k<q^{n-2}$, then $\mathcal{P}^{k}(u_{m})/u_{m}\in R_{n-2}$ and
$u_{m}^{\ell}\mathcal{P}^{k}(u_{m})\mathcal{P}^{i-k}(d_{j,m})=(\mathcal{P}^{k}(u_{m})/u_{m})u_{m}^{\ell+1}\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$
with $\xi_{n-1}$-degree at most $\ell+1$, as required. Suppose $q^{n-2}\leq k$. Then $\ell^{\prime}<\ell$ and $u_{m}^{\ell}\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell$. Since $\mathcal{P}^{k}(u_{m})\in R_{n-1}$ with with $\xi_{n-1}$-degree at most $1$, $u_{m}^{\ell}\mathcal{P}^{k}(u_{m})\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$, as required.
Suppose $i=(q-1)q^{n-2}+r$ with $r<q^{n-2}$ and $i\not\equiv_{(q)}0$. We know that $u_{m}^{q}\mathcal{P}^{i}(d_{j,m})\in R_{n-1}$ with with $\xi_{n-1}$-degree at most $q$. We will show that the $\xi_{n-1}$-degree at most $q-1$. Using the Cartan identity
@@
Since $u_{m}^{q}d_{j,m}\in R_{n-1}$ with $\xi_{n-1}$-degree $1$, we see that $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})\in R_{n-1}$ with with $\xi_{n-1}$-degree at most $q$. However, the coefficient of $\xi_{n-1}$ in $u_{m}^{q}d_{j,m}$ is given by $u_{m}^{q-1}(u_{m-1}d_{j-1,m-1})^{q}$, which means that the coefficient of $\xi_{n-1}\xi_{n-2}^{q-1}$ in $u_{m}^{q}d_{j,m}$ is $(u_{m-1}^{q}d_{j-1,m-1})^{q}$. Therefore the coefficient of $\xi_{n-1}^{q}$ in $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})$ is given by $\mathcal{P}^{r}((u_{m-1}^{q}d_{j-1,m-1})^{q})$. Since $r\not\equiv_{(q)}0$, $\mathcal{P}^{r}((u_{m-1}^{q}d_{j-1,m-1})^{q})=0$ and the $\xi_{n-1}$-degree of $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})$ is at most $q-1$. Since $i<q^{n-1}$ and $qk<i$, we have $k<q^{n-2}$. Thus $\mathcal{P}^{k}(u_{m})/u_{m}\in R_{n-2}$ and
@@
Write $r=r_{0}+qr^{\prime}$ with $0<r_{0}<q$. If $r^{\prime}=0$, then $i-kp<(q-1)q^{n-2}$ and $u_{m}^{q}\mathcal{P}^{i-qk}(d_{j,m})$ has $\xi_{n-1}$-degree at most $q-1$. If $r^{\prime}>0$, then the result follows by induction on $r^{\prime}$.
@@
$\mathcal{P}^{q^{n-3}}(c_{n-3-i,m-1}d_{i,m-1})=\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1})d_{i,m-1}^{q}+c_{n-3-i,m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}$
@@
$\xi_{n-2}$ $=$ $\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1}))d_{i,m-1}^{q}\right)$
$+\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\right)-\mathcal{P}^{q^{n-3}}(c_{n-3,m-1}).$
@@
$u_{m-1}^{q}\xi_{n-2}$ $=$ $\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1}))(u_{m-1}d_{i,m-1})^{q}\right)$
$+\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}u_{m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\right)$
$-u_{m-1}^{q}\mathcal{P}^{q^{n-3}}(c_{n-3,m-1}).$
By induction $u_{m-1}d_{i,m-1}\in R_{n-3}$ with $\xi_{n-3}$-degree $1$ and, since $a_{i}=q^{n-3}-q^{n-3-i}-1$, $u_{m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\in R_{n-3}$ with $\xi_{n-3}$-degree at most $q-1$. Therefore the right hand side of the relation is an element of $R_{n-3}$ and the relation lifts to give an element of $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}\subset S_{m}^{G_{m}}$. Comparing lead terms gives
$u_{m}$ $=$ $u_{m-1}^{q}\xi_{n-2}-\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1}))(u_{m-1}d_{i,m-1})^{q}\right)$
$-\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}u_{m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\right)$
$+u_{m-1}^{q}\mathcal{P}^{q^{n-3}}(c_{n-3,m-1}).$
Comparing the two expressions for $u_{m}$ and using the fact that $u_{m-1}^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})$ has $\xi_{n-3}$-degree at most $q-1$ while $(u_{m-1}d_{i,m-1})^{q}$ has $\xi_{n-3}$-degree $q$, we conclude that
$u_{m}=(\xi_{n-2}+c_{n-2,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(u_{m-1}d_{i,m-1})^{q}$
@@
$c_{n-2,m}=\mathcal{P}^{q^{n-3}}(c_{n-3,m-1})+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})\in R_{n-3}$
and $c_{n-2-i,m}=\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1})\in R_{n-3-i}$ for $0<i<m$. This completes the proof of part (a). ∎
7.5. Proof of Part (b)
The proof is by induction on $m$. For $m=2$, the result follows from Lemmas 6.1 and 6.4.
Suppose $m>2$. By induction, $u_{m-1}=M(0,m-1)+\delta_{0,m-1}$ and $u_{m-1}d_{i,m-1}=M(i,m-1)+\delta_{i,m-1}$. Substituting into the expression for $u_{m}$ from part (a) gives
$u_{m}$ $=$ $(\xi_{n-2}+c_{n-2,m})(M(0,m-1)+\delta_{0,m-1})^{q}$
$+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(M(i,m-1)+\delta_{i,m-1})^{q}$
$=$ $\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}M(i,m-1)^{q}+\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}\delta_{i,m-1}^{q}$
$+\sum_{i=0}^{m-1}(-1)^{i}c_{n-2-i}(M(i,m-1)^{q}+\delta_{i,m-1}^{q}).$
Observe that $M(0,m)=\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}M(i,m-1)^{q}$ and define
$\delta_{0,m}=\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}\delta_{i,m-1}^{q}+\sum_{i=0}^{m-1}(-1)^{i}c_{n-2-i}(M(i,m-1)^{q}+\delta_{i,m-1}^{q}).$
Then $u_{m}=M(0,m)+\delta_{0,m}$ with $\delta_{0,m}\in R_{n-2}$. Since $\nu(c_{n-2-i})>1$, $\nu(M(i,m-1))=1+q+\cdots+q^{m-2}$, and $\nu(\delta_{i,m-1})>1+q+\cdots+q^{m-2}$, we have $\nu(\delta_{0,m})>1+q+\cdots+q^{m-1}$.
Apply $\mathcal{P}^{e(i,m)}$ to $u_{m}$ to get $u_{m}d_{i,m}=\mathcal{P}^{e(i,m)}(M(0,m))+\mathcal{P}^{e(i,m)}(\delta_{0,m})$. Define $\delta_{i,m}:=u_{m}d_{i,m}-M(i,m)$ so that $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$, and observe that $\delta_{i,m}\in R_{n-1}$. Using Lemma 4.3,
@@ 7.5. Proof of Part (b)
It is clear that using the weighted grevlex order on $R_{n-1}$,
$\operatorname{LT}(M(0,m))=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1}+q^{m-2}+\cdots+1}$
@@
$\operatorname{LT}(M(i,m))=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1-i}+\cdots+1}\xi_{m}^{q^{m-1}+\cdots+q^{m-i}}.$
From Lemma 4.4, $\operatorname{LT}(\delta_{i,m})<\operatorname{LT}(M(i,m))$ for $0\leq i\leq m$. Therefore $\operatorname{LT}(u_{m})=\operatorname{LT}(M(0,m))$ and $\operatorname{LT}(u_{m}d_{im})=\operatorname{LT}(M(i,m))$, completing the proof part (b). ∎
7.6. Proof of Part (e)
The proof is by induction on $m$. For $m=2$, Equation 3 gives
$\xi_{2}^{q}=\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}+\xi_{0}^{q+1}r_{2,1}$
with $r_{2,1}\in R_{2}$. Define $\gamma_{1,2}^{(1)}=\gamma_{0,2}^{(1)}=0$ and $\gamma_{2,2}^{(1)}=-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{q+1}r_{2,1}$. Since $\nu(\gamma_{2,2}^{(1)})>q+1$, the result holds for $m=2$.
Suppose $m>2$. By induction we have parts (d) and (e) of the theorem for $S_{m-1}^{G_{m-1}}$. Therefore, for $1\leq i\leq m$, we have the relation
$0=u_{m-1}\left(\xi_{n-2-i}^{q^{i-1}}+\gamma_{n-2-i,m-1}^{(i-1)}-\sum_{j=1}^{m-1}(-1)^{j+1}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})d_{j,m-1}\right)$
in $R_{n-3}$. Define
$\gamma:=\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}u_{m}d_{j+1,m}.$
By part (a), $u_{m}d_{j,m}-(u_{m-1}d_{j-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$ for $1\leq j\leq m$, using the convention $d_{0,m-1}=1$. Therefore $\gamma\in R_{n-1}$. Furthermore $\gamma$ is linear in $\xi_{n-1}$ with coefficient
$\left(-\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})u_{m-1}d_{j,m-1}\right)^{q}=0.$
Hence $\gamma\in R_{n-2}$. From part (a), $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$. Since $u_{m}$ divides $\gamma$ in $S_{m}^{G_{m}}$ and $\gamma$ lies in $R_{n-2}$, we see that $\gamma\in u_{m}R_{n-2}$. Thus $\gamma/u_{m}\in R_{n-2}$.
Define an $(m+1)\times(m+1)$ matrix in $R_{n-1}$ by
$M_{m}^{(i)}:=\begin{pmatrix}\xi_{n-1-i}^{q^{i}}&\cdots&\xi_{m-1-i}^{q^{i}}\\ &M_{m}&\end{pmatrix}.$

### v3

@@ 7.2. Proof of Parts (a), (b) and (c)
For $i<q^{n-1}$, write $i=\ell q^{n-2}+r$ with $\ell<q$ and $r<q^{n-2}$. We will prove by induction on $i$ that $u_{m}^{\ell+1}\mathcal{P}^{i}(d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $\ell+1$. For $i=0$, we have $\ell=0$ and $u_{m}\mathcal{P}^{0}(d_{j,m})=u_{m}d_{j,m}\in R_{n}$ with $\xi_{n}$-degree $1$. For $0<i<q^{m-1}$, we have $u_{m}\mathcal{P}^{i}(d_{j,m})=\mathcal{P}^{i}(u_{m}d_{j,m})\in R_{n-1}$, giving $\mathcal{P}^{i}(d_{j,m})\in R_{n-1}$ (see Remark 4.7). For $i\geq q^{m-1}$, using the Cartan identity gives
@@
Since $u_{m}d_{j,m}\in R_{n}$ with $\xi_{n}$-degree $1$, we see that $\mathcal{P}^{i}(u_{m}d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $\ell+1$. For $k>0$, write $i-k=\ell^{\prime}q^{n-2}+r^{\prime}$ with $\ell^{\prime}\leq\ell$ and $r^{\prime}<q^{n-2}$. Since ik<$i$-k<i, the induction hypothesis gives $u_{m}^{\ell^{\prime}+1}\mathcal{P}^{i-k}(d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $\ell^{\prime}+1$. If $k<q^{n-2}$, then $\mathcal{P}^{k}(u_{m})/u_{m}\in R_{n-1}$ and
$u_{m}^{\ell}\mathcal{P}^{k}(u_{m})\mathcal{P}^{i-k}(d_{j,m})=(\mathcal{P}^{k}(u_{m})/u_{m})u_{m}^{\ell+1}\mathcal{P}^{i-k}(d_{j,m})\in R_{n}$
with $\xi_{n}$-degree at most $\ell+1$, as required. Suppose $q^{n-2}\leq k$. Then $\ell^{\prime}<\ell$ and $u_{m}^{\ell}\mathcal{P}^{i-k}(d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $\ell$. Since $\mathcal{P}^{k}(u_{m})\in R_{n}$ with $\xi_{n}$-degree at most $1$, $u_{m}^{\ell}\mathcal{P}^{k}(u_{m})\mathcal{P}^{i-k}(d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $\ell+1$, as required.
Suppose $i=(q-1)q^{n-2}+r$ with $r<q^{n-2}$ and $i\not\equiv_{(q)}0$. We know that $u_{m}^{q}\mathcal{P}^{i}(d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $q$. We will show that the $\xi_{n}$-degree at most q1q-1. Using the Cartan identity
@@
Since $u_{m}^{q}d_{j,m}\in R_{n}$ with $\xi_{n}$-degree $1$, we see that $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $q$. However, the coefficient of $\xi_{n}$ in $u_{m}^{q}d_{j,m}$ is given by $u_{m}^{q-1}(u_{m-1}d_{j-1,m-1})^{q}$, which means that the coefficient of $\xi_{n}\xi_{n-1}^{q-1}$ in $u_{m}^{q}d_{j,m}$ is $(u_{m-1}^{q}d_{j-1,m-1})^{q}$. Therefore we have that the coefficient of $\xi_{n}^{q}$ in $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})$ is given by $\mathcal{P}^{r}((u_{m-1}^{q}d_{j-1,m-1})^{q})$. Since $r\not\equiv_{(q)}0$, $\mathcal{P}^{r}((u_{m-1}^{q}d_{j-1,m-1})^{q})=0$ and the $\xi_{n}$-degree of $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})$ is at most q1q-1. Since $i<q^{n-1}$ and $qk<i$, we have $k<q^{n-2}$. Thus $\mathcal{P}^{k}(u_{m})/u_{m}\in R_{n-1}$ and
@@
Write $r=r_{0}+qr^{\prime}$ with $0<r_{0}<q$. If $r^{\prime}=0$, then $i-kq<(q-1)q^{n-2}$ and $u_{m}^{q}\mathcal{P}^{i-qk}(d_{j,m})$ has $\xi_{n}$-degree at most q1q-1. If $r^{\prime}>0$, then the result follows by induction on $r^{\prime}$.
@@
$\mathcal{P}^{q^{n-3}}(c_{n-2-i,m-1}d_{i,m-1})=\mathcal{P}^{q^{n-3-i}}(c_{n-2-i,m-1})d_{i,m-1}^{q}+c_{n-2-i,m-1}^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})$
@@
$\xi_{n-1}$ $=$ $\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-1-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-2-i,m-1}))d_{i,m-1}^{q}\right)$
$+\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+c_{n-2-i,m-1})^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})\right)-\mathcal{P}^{q^{n-3}}(c_{n-2,m-1}).$
@@
$u_{m-1}^{q}\xi_{n-1}$ $=$ $\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-1-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-2-i,m-1}))(u_{m-1}d_{i,m-1})^{q}\right)$
$+\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+c_{n-2-i,m-1})^{q}u_{m-1}^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})\right)$
$-u_{m-1}^{q}\mathcal{P}^{q^{n-3}}(c_{n-2,m-1}).$
By induction $u_{m-1}d_{i,m-1}\in R_{n-2}$ with $\xi_{n-2}$-degree $1$ and, since $a_{i}=q^{n-3}-q^{n-3-i}-1$, $u_{m-1}^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})\in R_{n-2}$ with $\xi_{n-2}$-degree at most q1q-1. Therefore the right hand side of the relation is an element of $R_{n-2}$ and the relation lifts to give an element of $\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}\subset S_{m}^{G_{m}}$. Comparing the coefficient of $\xi_{n-1}$ gives
$u_{m}$ $=$ $u_{m-1}^{q}\xi_{n-1}-\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-1-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-2-i,m-1}))(u_{m-1}d_{i,m-1})^{q}\right)$
$-\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+c_{n-2-i,m-1})^{q}u_{m-1}^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})\right)$
$+u_{m-1}^{q}\mathcal{P}^{q^{n-3}}(c_{n-2,m-1}).$
Comparing this expression for $u_{m}$ with the expression coming from Equation 7 and using the fact that $u_{m-1}^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})$ has $\xi_{n-2}$-degree at most q1q-1 while $(u_{m-1}d_{i,m-1})^{q}$ has $\xi_{n-2}$-degree $q$, we conclude that
$u_{m}=(\xi_{n-1}+c_{n-1,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-1-i}+c_{n-1-i,m})(u_{m-1}d_{i,m-1})^{q}$
@@
$c_{n-1,m}=\mathcal{P}^{q^{n-3}}(c_{n-2,m-1})+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m-1})^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})\in R_{n-2}$
and $c_{n-1-i,m}=\mathcal{P}^{q^{n-3-i}}(c_{n-2-i,m-1})\in R_{n-2-i}$ for $0<i<m$. This completes the proof of parts (a), (b) and (c). ∎
7.3. Proof of Part (d)
The proof is by induction on $m$. For $m=2$, the result follows from Lemmas 6.1 and 6.6.
Suppose $m>2$. By induction, $u_{m-1}=M(0,m-1)+\delta_{0,m-1}$ and $u_{m-1}d_{i,m-1}=M(i,m-1)+\delta_{i,m-1}$. Substituting into the expression for $u_{m}$ from part (c) gives
$u_{m}$ $=$ $(\xi_{n-1}+c_{n-1,m})(M(0,m-1)+\delta_{0,m-1})^{q}$
$+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-1-i}+c_{n-1-i,m})(M(i,m-1)+\delta_{i,m-1})^{q}$
$=$ $\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-1-i}M(i,m-1)^{q}+\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-1-i}\delta_{i,m-1}^{q}$
$+\sum_{i=0}^{m-1}(-1)^{i}c_{n-1-i,m}(M(i,m-1)^{q}+\delta_{i,m-1}^{q}).$
Observe that $M(0,m)=\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-1-i}M(i,m-1)^{q}$ and define
$\delta_{0,m}=\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-1-i}\delta_{i,m-1}^{q}+\sum_{i=0}^{m-1}(-1)^{i}c_{n-1-i,m}(M(i,m-1)^{q}+\delta_{i,m-1}^{q}).$
Then $u_{m}=M(0,m)+\delta_{0,m}$ with $\delta_{0,m}\in R_{n-1}$. Since $c_{j,m}\in R_{j-1}$ and ${S\mbox{-}\deg}(c_{j,m})={S\mbox{-}\deg}(\xi_{j})=q^{j-1}+1$, we see that $\nu(c_{j,m})>1$. Hence $\nu(M(i,m-1))=1+q+\cdots+q^{m-2}$ and $\nu(\delta_{i,m-1})>1+q+\cdots+q^{m-2}$, we have $\nu(\delta_{0,m})>1+q+\cdots+q^{m-1}$.
Apply $\mathcal{P}^{e(i,m)}$ to $u_{m}$ to get $u_{m}d_{i,m}=\mathcal{P}^{e(i,m)}(M(0,m))+\mathcal{P}^{e(i,m)}(\delta_{0,m})$. Define $\delta_{i,m}:=u_{m}d_{i,m}-M(i,m)$ so that $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$, and observe that $\delta_{i,m}\in R_{n}$. Using Lemma 4.3,
@@ 7.3. Proof of Part (d)
It is clear that using the weighted revlex order on $R_{n}$,
$\operatorname{LT}(M(0,m))=(-1)^{\lfloor m/2\rfloor}\xi_{m}^{q^{m-1}+q^{m-2}+\cdots+1}$
@@
$\operatorname{LT}(M(i,m))=(-1)^{\lfloor m/2\rfloor}\xi_{m}^{q^{m-1-i}+\cdots+1}\xi_{m+1}^{q^{m-1}+\cdots+q^{m-i}}.$
From Lemma 4.4, $\operatorname{LT}(\delta_{i,m})<\operatorname{LT}(M(i,m))$ for $0\leq i\leq m$. Therefore $\operatorname{LT}(u_{m})=\operatorname{LT}(M(0,m))$ and $\operatorname{LT}(u_{m}d_{im})=\operatorname{LT}(M(i,m))$, completing the proof part (d). ∎
7.4. Proof of Part (f)
The proof is by induction on $m$. The base case, $m=2$, is given in Section 6. Suppose $m>2$. By induction we have parts (e) and (f) of the lemma for $S_{m-1}^{G_{m-1}}$. Therefore, for $1\leq i\leq m-1$, we have the relation
$0=u_{m-1}\left(\xi_{n-1-i}^{q^{i-1}}+\gamma_{n-1-i,m-1}^{(i-1)}-\sum_{j=1}^{m-1}(-1)^{j+1}(\xi_{n-1-i-j}^{q^{i-1}}+\gamma_{n-1-i-j,m-1}^{(i-1)})d_{j,m-1}\right)$
in $R_{n-2}$. Define
$\gamma^{(i)}:=\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-1-i-j}^{q^{i-1}}+\gamma_{n-1-i-j,m-1}^{(i-1)})^{q}u_{m}d_{j+1,m}.$
By part (c), $u_{m}d_{j,m}-(u_{m-1}d_{j-1,m-1})^{q}\xi_{n}\in R_{n-1}$ for $1\leq j\leq m$, using the convention $d_{0,m-1}=1$. Therefore $\gamma^{(i)}\in R_{n}$. Furthermore $\gamma^{(i)}$ is linear in $\xi_{n}$ with coefficient
$\left(-\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-1-i-j}^{q^{i-1}}+\gamma_{n-1-i-j,m-1}^{(i-1)})u_{m-1}d_{j,m-1}\right)^{q}=0.$
Hence $\gamma^{(i)}\in R_{n-1}$. From part (a), $\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}$. Since $u_{m}$ divides $\gamma^{(i)}$ in $S_{m}^{G_{m}}$ and $\gamma^{(i)}$ lies in $R_{n-1}$, we see that $\gamma^{(i)}\in u_{m}R_{n-1}$ (see Remark 4.7). Thus $\gamma^{(i)}/u_{m}\in R_{n-1}$.
Define an $(m+1)\times(m+1)$ matrix over $R_{n}$ by
$M_{m}^{(i)}:=\begin{pmatrix}\xi_{n-i}^{q^{i}}&\cdots&\xi_{m-i}^{q^{i}}\\ &M_{m}&\end{pmatrix}$.